# PiKV: KV Cache Management System for Mixture of Experts

### v1

@@
Abstract.
As large language models continue to scale up in both size and context length, the memory and communication cost of key-value (KV) cache storage has become a major bottleneck in multi-GPU and multi-node inference. While MoE-based architectures sparsify computation across experts, the corresponding KV caches remain dense and globally synchronized, resulting in significant overhead.
@@ Abstract.
PiKV is recently publicly available as an open-source software library: https://github.com/NoakLiu/PiKV. Experiments details is recorded at: https://github.com/NoakLiu/PiKV/Experimental_Results. We also have PiKV integrated with Nvidia kvpress for acceleration, details see https://github.com/NoakLiu/PiKVpress. PiKV is still a living project, aiming to become a comprehesive KV Cache management system for MoE Architectures.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/NoakLiu/PiKV.
1. Introduction
@@
However, serving such models introduces significant system-level challenges. During inference, each token generation requires attending to the entire KV cache from prior tokens. For a 7B-scale MoE model with 128K context and 16 experts, the full KV cache can occupy $>$24GB of memory and incur excessive communication latency across GPUs and nodes. Even with FlashAttention-style optimizations (Dao et al., 2022), the need to load and attend to dense KV structures becomes the dominant bottleneck, especially in autoregressive decoding.
@@
Figure 1. PiKV Framework
Figure 2. KV cache memory usage comparison. Left: absolute memory usage of different methods; Right: memory reduction percentage compared to Full KV cache.
Figure 3. Latency performance comparison. Left: latency at different sequence lengths; Right: speedup ratio of PiKV compared to other methods.
We propose PiKV, a parallel distributed KV caching system tailored for sparse mixture of expert models training and inference. As shown in Figure 1, PiKV includes three synergistic components:
[…]
Our experimental results demonstrate significant improvements in both memory efficiency and inference latency. As shown in Figures 2 and 3, PiKV achieves up to 3.9$\times$ memory reduction and 1.7$\times$ latency improvement compared to full KV cache approaches, while maintaining competitive accuracy across various benchmarks.
@@
2. Preliminary
2.1. Sparse MoE Inference Meets Long-Context Bottlenecks
The rise of long-context language models has transformed downstream tasks such as multi-document reasoning, conversational memory, and retrieval-augmented generation. However, inference with context lengths beyond 32K tokens introduces prohibitive memory and latency overhead, especially when deployed with Mixture-of-Experts (MoE) models (Lepikhin et al., 2020; Fedus et al., 2022), which already impose routing and communication complexities. Despite the potential for sparse computation, current MoE systems suffer from dense KV cache replication, non-adaptive expert selection, and cache scheduling agnostic to query dynamics.
[…]
Challenge 1: Expert-Sharded KV Fragmentation. Token-level routing naturally shards the KV cache across devices. However, naive per-expert partitioning breaks temporal locality and introduces fragmented access patterns:
$\text{KV}_{t}^{(e)}\in\mathbb{R}^{k\times d},\quad\text{for }e\in\mathcal{R}(q_{t})$
where $\mathcal{R}$ is the router assigning experts to token $t$. Efficient access thus requires query-aware retrieval across heterogeneous and often incomplete cache segments.
Challenge 2: Latency Bottleneck from Sparse Lookup. Sparse MoE inference reduces compute, but not necessarily latency: each query $q_{t}$ must attend over $k$ expert KV sets, often spread across GPUs. Without localized cache routing and pre-filtering, the expected latency is:
$\mathbb{E}[\text{Latency}]\sim\mathcal{O}(k\cdot T_{\text{lookup}}+T_{\text{sync}})$
where $T_{\text{sync}}$ arises from inter-GPU communication.
Challenge 3: Non-coordinated Routing, Compression, and Scheduling. Prior systems treat expert routing (Du et al., 2022), KV compression (Liu et al., 2025; Liu, 2024), and cache scheduling (Xiao et al., 2024; Tang et al., 2024) as disjoint modules. This results in inconsistent memory policies: a router might route to an expert whose KV cache has already evicted the relevant tokens.
PiKV bridges this gap. We introduce a unified framework where:
[…]
  • Compression is hierarchical and expert-partitioned, minimizing redundancy without affecting reuse;
  • Scheduling is jointly optimized with routing, based on token-level saliency and inter-expert redundancy.
By reframing the inference stack around the KV cache as a central abstraction, PiKV orchestrates routing, compression, and eviction as a coupled optimization problem:
$\min_{\mathcal{R},\mathcal{C},\mathcal{S}}\quad\mathbb{E}_{q\sim Q}\left[\text{Latency}(q)+\lambda_{1}\cdot\text{Memory}(q)-\lambda_{2}\cdot\text{Fidelity}(q)\right]$
subject to:
$\begin{cases}\mathcal{R}(q)\subseteq\{1,\dots,E\},&\text{expert routing}\\ \mathcal{C}:\text{KV}\to\text{CompressedKV},&\text{KV compression}\\ \mathcal{S}:\text{Cache}\to\text{EvictableSet},&\text{cache scheduling}\end{cases}$
This coupled optimization leads to concrete system benefits, including lower latency, reduced bandwidth, and improved cache hit rates in long-context MoE inference, as demonstrated in our experiments.
2.2. Cache Fragmentation and Memory Contention in Sparse MoE Inference
While sparse MoE architectures reduce FLOPs by activating a limited set of experts per token (Du et al., 2022; Shazeer et al., 2017), they do not inherently solve the challenges of caching under long-context settings. In practical deployments, sparse compute paths are coupled with fragmented KV cache access, cross-device lookups, and incoherent compression-eviction policies.
[…]
$\text{Attn}(x_{t})=\sum_{e\in\mathcal{R}(x_{t})}\sum_{\tau<t}\alpha_{\tau}^{(e)}\cdot\text{value}_{\tau}^{(e)}$
However, in expert-sharded memory layouts, the retrieval:
$\text{KV}^{(e)}_{t}=\texttt{Retrieve}(x_{<t},\mathcal{K}^{(e)})$
may involve only partial or stale KV views due to eviction or incomplete synchronization, leading to degraded attention fidelity.
Challenge 1: Fragmentation from Routing-Induced Sharding. Token-level routing scatters KV tokens across experts and devices. Temporal prefixes for a given token may reside in different expert-local caches, breaking attention locality.
Challenge 2: Latency from Distributed Lookups. Each token’s attention involves multiple inter-expert and possibly inter-device KV lookups:
$\text{Latency}_{\text{token}}\sim\sum_{e\in\mathcal{R}(x_{t})}\left(T_{\text{fetch}}^{(e)}+T_{\text{decode}}^{(e)}\right)$
which becomes the bottleneck in low-latency scenarios.
Challenge 3: Misalignment between Routing, Compression, and Scheduling. A token might be routed to an expert whose recent KV has been heavily compressed or evicted. This cache-state mismatch leads to performance degradation that routing alone cannot mitigate.
PiKV addresses these deployment issues with a cache-centric pipeline.
  • Routing is KV-aware: Tokens avoid experts with recently flushed or compressed caches.
  • Compression is hierarchical and page-based: Using structures like PyramidKV and ChunkKV, we compress without fragmenting token boundaries.
  • Scheduling is query-sensitive: We compute token-level reuse scores to inform eviction priorities under constrained GPU memory.
The final optimization problem becomes:
$\min_{\mathcal{R},\mathcal{S},\mathcal{C}}\;\mathbb{E}_{x_{t}}\left[\sum_{e\in\mathcal{R}(x_{t})}\left(T_{\text{fetch}}^{(e)}+\lambda\cdot\text{MemCost}^{(e)}-\mu\cdot\text{HitRate}^{(e)}\right)\right]$
subject to:
$\begin{cases}\mathcal{R}(x_{t})\subseteq\{1,\dots,E\},&\text{top-$k$ routing constraint}\\ \mathcal{S}:\text{Cache}\rightarrow\text{EvictableSet},&\text{stream-aware scheduling}\\ \mathcal{C}:\mathcal{K}^{(e)}\rightarrow\mathcal{K}_{\text{compressed}}^{(e)}&\text{per-expert compression}\end{cases}$
PiKV thus bridges high-level routing policy with low-level cache mechanics, enabling scalable sparse inference over long context contexts with high cache hit rate and small latency.
2.3. MoE Compression Meets Streaming: Coordinating Memory Pressure with Fidelity Guarantees
KV cache compression and scheduling are critical to long-context inference, especially under constrained memory and bandwidth. Prior works propose numerous token pruning and quantization strategies—yet few are optimized for sparse expert inference where cache locality, query adaptivity, and temporal reuse must be jointly considered.
Concretely, in autoregressive decoding, each token $x_{t}$ must attend to a subset of previous key-value entries:
$\text{Attn}(x_{t})=\sum_{\tau<t}\alpha_{\tau}\cdot V_{\tau},\quad\text{where }(K_{\tau},V_{\tau})\in\texttt{KVCache}$
Given a maximum memory budget $M$, the system must dynamically decide which past entries to compress, evict, or retain, under constraints of attention recall:
$\text{KVCache}_{t}=\mathcal{C}_{t}\circ\mathcal{S}_{t}(\text{KVCache}_{t-1},x_{t})$
where $\mathcal{C}_{t}$ is a compression operator (e.g., quantization, low-rank projection) and $\mathcal{S}_{t}$ is a scheduling rule for eviction or retention.
Challenge 1: Tradeoff between Fidelity and Throughput. Aggressive pruning improves throughput but risks degrading output quality. Let $\hat{V}_{\tau}=\mathcal{C}(V_{\tau})$ be the compressed value. The fidelity loss is:
$\mathcal{L}_{\text{fidelity}}=\sum_{\tau<t}\|\alpha_{\tau}V_{\tau}-\alpha_{\tau}\hat{V}_{\tau}\|_{2}^{2}$
Minimizing this while staying under memory budget leads to a non-trivial scheduling-compression problem.
Challenge 2: Streaming Context Requires On-the-Fly Selection. Unlike fixed-length prompts, many applications (e.g., chat history, online document generation) involve unbounded inputs. Systems like StreamingLLM (Xiao et al., 2024) and Quest (Tang et al., 2024) propose scoring-based methods for token eviction. However, they assume centralized scoring access and do not generalize well to expert-sharded caches.
Challenge 3: Expert-local Compression Conflicts with Global Utility. Compression is often applied uniformly, ignoring how useful a KV entry is to downstream experts. Ideally, compression should be informed by:
$u(e,\tau)=\mathbb{E}_{x_{t}\in\mathcal{R}^{-1}(e)}\left[\alpha_{\tau}^{(t)}\right]$
where $u(e,\tau)$ denotes the expected utility of token $\tau$ for expert $e$—a measure that bridges compression and routing.
PiKV resolves this via hierarchical, token-activity-aware compression and scheduling.
  • Compression is modular and stratified. PiKV integrates multiple strategies—low-rank approximation (e.g., LoRA (Hu et al., 2022)), multi-resolution clustering (e.g., PyramidKV (Cai et al., 2024)), and chunk-level merging (e.g., ChunkKV)—into a layered framework:
    $\mathcal{C}=\mathcal{C}^{\text{rank}}\circ\mathcal{C}^{\text{quant}}\circ\mathcal{C}^{\text{chunk}}$
  • Scheduling is stream-sensitive and reusability-aware. We define a scoring function:
    $s_{\tau}=\texttt{reuse}(\tau)+\gamma\cdot\texttt{similarity}(\tau,x_{t})$
    where $\texttt{reuse}(\tau)$ tracks future access likelihood and similarity measures token-query match via cosine score in KV space.
  • Eviction is batched and context-aware. Instead of per-token eviction, PiKV evicts token groups based on redundancy within expert-local caches.
Finally, the compression-scheduling tradeoff is formally captured by:
$\min_{\mathcal{C},\mathcal{S}}\;\mathbb{E}_{x_{t}}\left[\mathcal{L}_{\text{fidelity}}(x_{t})+\beta\cdot\mathcal{L}_{\text{throughput}}(x_{t})\right]\quad\text{s.t. }\|\text{KVCache}_{t}\|\leq M$
By explicitly modeling token reuse, activity, and expert overlap, PiKV’s compression and scheduling engine achieves high cache hit rate with bounded memory and negligible generation degradation—as shown across long-context benchmarks in our experiments.
3. Methodology
@@
  • Sparsity-aware serving: Only a small set of experts and KV pages are relevant per query;
  • Resource-constrained scheduling: The memory and bandwidth budget must be dynamically partitioned across queries, experts, and streams.
@@
All components are executed in an asynchronous pipeline orchestrated by a general decoding loop, as shown in Algorithm 1. Each submodule operates independently but passes metadata to adjacent stages to inform decisions. The comprehensive system architecture is illustrated in Figure 1, while detailed ablation studies demonstrating the contribution of each component are presented in Figure 4.
@@
3.1. PiKV Expert-Sharded Storage
@@ 3.2. PiKV Routing
$\mathcal{R}_{\textbf{LB}}$ TopK + load balance $-\alpha(\mu_{e}-\bar{\mu})$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{P}}$ Cache-aware (PiKVRouter) $-\lambda\log\!\bigl{(}1+\text{miss}_{e}\bigr{)}$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{E}}$ Entropy-penalised LB (EPLB) $-\beta\,H(p_{e})$ $\mathcal{O}(E)$
@@
Table 1. PiKV routing methods and their computational profiles ($E$ = experts).
@@
Notations and Mathematical Framework: Let $d$ be the hidden size, $h$ be the head width, $E$ be the total expert count, $k\ll E$ be the active experts per token, $B$ be the batch size, and $L$ be the sequence length. We denote the hidden state as $h\in\mathbb{R}^{d}$, the attention matrix as $H\in\mathbb{R}^{d\times L}$, and the expert routing weights as $r\in\mathbb{R}^{E}$.
For dense attention, we access all experts per token, while for sparse routing with PiKV, only $k$ experts are activated. We define the router logits as $r^{\prime}=W_{r}\cdot h$, where $W_{r}\in\mathbb{R}^{E\times d}$ is the routing matrix.
Memory Traffic Analysis of PiKV Routing.
Memory I/O Formulation: Following the MoE evaluation methodology, we define the memory I/O operations for Key-Value cache access. The memory traffic can be decomposed into three components: (1) Key matrix access, (2) Value matrix access, and (3) attention computation overhead.
For dense attention, the total memory I/O is:
[…]
where the first term represents Key-Value matrix access and the second term accounts for attention computation overhead.
For sparse routing with PiKV:
[…]
Cache Performance and Locality Analysis: We define reuse distance as a measure of temporal locality in cache access patterns. For dense attention:
$\operatorname{RD}_{\text{dense}}=\frac{L}{E}$
For sparse routing with PiKV:
$\operatorname{RD}_{\text{sparse}}=\frac{L}{k}$
The cache hit rate can be approximated using the reuse distance ratio:
$\operatorname{RD}_{\text{dense}}\!=\!\frac{L}{E},\qquad\operatorname{RD}_{\text{sparse}}\!=\!\frac{L}{k}\;\;\Longrightarrow\;\;\text{cache hit-rate}\approx\frac{k}{E}.$
Arithmetic Intensity and Hardware Utilization: Following the MoE evaluation methodology, we define arithmetic intensity (IN) as the ratio of computational operations to memory operations:
$\text{IN}_{\text{dense}}=\frac{\text{FLOPS}_{\text{dense}}}{\text{I/O}_{\text{dense}}}=\frac{B\,L\,h\,E}{2B\,L\,h\,E+B\,L\,d\,E}=\frac{h}{2h+d}$
$\text{IN}_{\text{sparse}}=\frac{\text{FLOPS}_{\text{sparse}}}{\text{I/O}_{\text{sparse}}}=\frac{B\,L\,h\,k}{2B\,L\,h\,k+B\,L\,d\,k}=\frac{h}{2h+d}$
While arithmetic intensity remains constant, the absolute memory traffic reduction enables larger effective batch sizes and improved hardware utilization.
Throughput Analysis Based on Roofline Model: Following the Roofline model approach, we can express the throughput improvement as:
$\text{Throughput}_{\text{sparse}}=\min\left(\text{Peak\_Compute},\frac{\text{Peak\_Memory}\times\text{AI}_{\text{sparse}}}{\text{I/O}_{\text{sparse}}}\right)$
The throughput scaling factor is:
[…]
This demonstrates that PiKV achieves linear throughput scaling with the expert reduction ratio, subject to hardware constraints.
Load Balancing and Expert Utilization: We define expert utilization efficiency as:
$\eta_{\text{util}}=\frac{k}{E}\times\frac{\text{Active\_Experts}}{\text{Total\_Experts}}$
For optimal load balancing, we require:
$\eta_{\text{util}}\geq\eta_{\text{threshold}}$
where $\eta_{\text{threshold}}$ is the minimum utilization threshold for efficient hardware usage.
3.3. PiKV Compression

### v2

# PiKV: KV Cache Management System for MoE Architecture

@@
Abstract
As large-scale language models continue to scale up in both size and context length, the memory and communication cost of key-value (KV) cache storage has become a major bottleneck in multi-GPU and multi-node inference. While MoE-based architectures sparsify computation across experts, the corresponding KV caches remain dense and globally synchronized, resulting in significant overhead.
@@ Abstract
PiKV is recently publicly available as an open-source software library: https://github.com/NoakLiu/PiKV. Experiments details is recorded at: Experimental Results. We also have PiKV integrated with Nvidia kvpress for acceleration, details see PiKVpress. PiKV is still a living project, aiming to become a comprehesive KV Cache management system for MoE Architectures.
Machine Learning, ICML
1 Introduction
@@
However, serving such models introduces significant system-level challenges. During inference, each token generation requires attending to the entire KV cache from prior tokens. For a 7B-scale MoE model with 128K context and 16 experts, the full KV cache can occupy >>24GB of memory and incur excessive communication latency across GPUs and nodes. Even with FlashAttention-style optimizations (Dao et al., 2022), the need to load and attend to dense KV structures becomes the dominant bottleneck, especially in autoregressive decoding.
@@
Figure 1: PiKV Framework
We propose PiKV, a parallel distributed KV caching system tailored for sparse mixture of expert models training and inference. As shown in Figure 1, PiKV includes three synergistic components: (1) an expert-sharded distributed KV cache layout across multi-GPU or multi-node compute, (2) a sparse expert routing layer that dynamically selects top-$k$ experts per query, and (3) an adaptive stream scheduler that uses activity-based eviction to retain only high-utility KV entries.
@@
2 Related Work
2.1 Long-context LLMs
With the rapid growth in model size and sequence length, long-context LLMs have gained significant attention. Architectures such as Transformer-XL (Dai et al., 2019), Longformer (Beltagy et al., 2020), and BigBird (Zaheer et al., 2020) explored architectural sparsity to improve scalability. In practice, most LLMs adopt Rotary Position Embedding (RoPE) (Su et al., 2024), which has been extended to longer contexts through rescaling (Chen et al., ). Notably, Yarn-LLaMA (Peng et al., 2023) expanded the 4K-token LLaMA-2 model to 32K and 128K tokens, respectively. Context length scaling techniques have pushed limits beyond 1M tokens (An et al., 2024). For long-document semantic understanding, hierarchical memory structures (Liu and Yu, 2025b) organize content into local segment graphs and a compact global summary, reducing worst-case semantic parsing complexity from $\mathcal{O}(N^{2})$ to near-linear. Serving these long-context models introduces substantial KV cache pressure and decoding latency. PiKV addresses this by combining distributed KV cache design with token-level adaptive routing and scheduling.
2.2 Sparse Expert Routing and KV Lookup
MoE-based models (Lepikhin et al., 2020; Fedus et al., 2022; Du et al., 2022) reduce compute cost by activating a small subset of experts per token. However, inference-time memory and KV access patterns remain dense in most implementations. DeepSeek-V2 (Liu et al., 2024) and M6-MoE (Zheng et al., 2023) explored dynamic expert selection and sparsity-aware routing. PiOne (Lu et al., 2024) and RingAttention (Liu et al., 2023) explored multi-GPU KV alignment but did not optimize for token-expert selectivity. At the hardware level, CXL-SpecKV (Liu and Yu, 2026) leverages CXL interconnects and FPGA accelerators to disaggregate KV caches to remote memory, combining speculative prefetching with on-chip compression to reduce GPU memory pressure in datacenter deployments. Our work differs by tightly coupling token-level sparse routing with expert-sharded KV cache layouts and a lightweight backend gating network, reducing lookup and communication cost simultaneously.
2.3 KV Compression and Stream-aware Scheduling
To alleviate memory and bandwidth bottlenecks, several methods have been proposed to compress and truncate KV caches. H2O (Zhang et al., 2023) prioritizes tokens with high cumulative attention weights. FastGen (Sheng et al., 2023) selects KV entries based on token type sensitivity. StreamingLLM (Xiao et al., 2024) introduce streaming sinks and proxy scoring methods to handle infinite-length texts. Quest (Tang et al., 2024) and TOVA (He et al., 2025) further propose query-aware KV selection by evaluating relevance to the current query.
TinyServe (Liu and Yu, 2025e) introduces query-aware page selection via bounding-box metadata to estimate per-block attention relevance, enabling selective KV loading with a fused CUDA kernel that integrates page scoring, sparse memory access, and masked attention in a single pass. LLMEasyQuant (Liu and Yu, 2025c) further reduces KV footprint through scalable quantization tailored for parallel and distributed inference. Cache acceleration has also been studied in diffusion transformer models: FastCache (Liu et al., 2025b) employs learnable linear approximation to skip redundant cache computations, while AdaCorrection (Liu et al., 2026) applies adaptive offset correction to maintain generation fidelity under aggressive caching—insights that transfer to autoregressive LLM caching pipelines.
While effective, most methods either discard unused tokens too early or require full cache for scoring. In contrast, PiKV retains a unified KV cache layout with on-demand query-aware page loading, hierarchical compression (e.g., LoRA (Hu et al., 2022), PyramidKV (Cai et al., 2024)), and streaming-aware eviction based on token activity scores. This provides a practical tradeoff between throughput, memory, and attention fidelity.
2.4 Efficient ML Systems and Adaptive Planning
Beyond KV cache management, broader system-level efforts have demonstrated that adaptive compute allocation improves efficiency across diverse ML workloads. GraphSnapShot (Liu and Yu, 2025a) accelerates graph machine learning by decoupling the sampling and feature aggregation stages, showing that principled system co-design yields significant throughput gains. EchoRL (Liu et al., 2025a) exploits experience replay to guide adaptive planning in reinforcement learning, illustrating how learned scheduling policies reduce wasted computation. Mt2St (Liu and Yu, 2025d) studies adaptive multi-task to single-task distillation, demonstrating that task-aware resource allocation can simplify complex inference pipelines—an insight directly applicable to expert-routing decisions in MoE serving. CoPrimeEEG (Yu et al., 2026) applies co-prime sub-Nyquist sampling with dual-branch reconstruction to EEG signals, illustrating how structured sparsity and adaptive reconstruction principles extend across data modalities. These works collectively motivate PiKV’s philosophy of treating cache management as a dynamic, query-driven decision problem rather than a static policy.
3 Methodology
@@
  • C1 (Sparsity-aware serving): Only a small set of experts and KV pages are relevant per query;
  • C2 (Resource-constrained scheduling): The memory and bandwidth budget must be dynamically partitioned across queries, experts, and streams.
@@
All components are executed in an asynchronous pipeline orchestrated by a general decoding loop, as shown in Algorithm 1. Each submodule operates independently but passes metadata to adjacent stages to inform decisions.
@@
3.1 PiKV Expert-Sharded Storage
@@ 3.2 PiKV Routing
Table 1: PiKV routing methods and their computational profiles ($E$ = experts).
@@
$\mathcal{R}_{\textbf{LB}}$ TopK + load balance $\alpha(\mu_{e}-\bar{\mu})$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{P}}$ Cache-aware (PiKVRouter) $\lambda\log\!\bigl(1+\text{miss}_{e}\bigr)$ $\mathcal{O}(E)$
$\mathcal{R}_{\textbf{E}}$ Entropy-penalised LB (EPLB) $\beta\,H(p_{e})$ $\mathcal{O}(E)$
@@
Let $B$ be batch size, $L$ sequence length, $h$ head width, $E$ experts, $k\!\ll\!E$ routed per token.
$\boxed{C_{\text{dense}}=B\,L\,h\,E},\qquad\boxed{C_{\text{sparse}}=B\,L\,h\,k},\quad\Longrightarrow\quad\text{speed\!-\!up}=\frac{E}{k}.$
Memory traffic (Key–Value fetch, bytes):
$M_{\text{dense}}=2B\,L\,h\,E,\qquad M_{\text{sparse}}=2B\,L\,h\,k,\quad\text{relief}=\frac{E}{k}.$
Reuse-distance:
$\operatorname{RD}_{\text{dense}}\!=\!\frac{L}{E},\qquad\operatorname{RD}_{\text{sparse}}\!=\!\frac{L}{k}\;\;\Longrightarrow\;\;\text{hit-rate}\approx\frac{k}{E}.$
3.3 PiKV Compression